### v1

# ChatReID: Open-ended Interactive Person Retrieval via Hierarchical Progressive Tuning for Vision Language Models

Abstract
Person re-identification (Re-ID) is a critical task in human-centric intelligent systems, enabling consistent identification of individuals across different camera views using multi-modal query information. Recent studies have successfully integrated LVLMs with person Re-ID, yielding promising results. However, existing LVLM-based methods face several limitations. They rely on extracting textual embeddings from fixed templates, which are used either as intermediate features for image representation or for prompt tuning in domain-specific tasks. Furthermore, they are unable to adopt the VQA inference format, significantly restricting their broader applicability. In this paper, we propose a novel, versatile, one-for-all person Re-ID framework, ChatReID. Our approach introduces a Hierarchical Progressive Tuning (HPT) strategy, which ensures fine-grained identity-level retrieval by progressively refining the model’s ability to distinguish pedestrian identities. Extensive experiments demonstrate that our approach outperforms SOTA methods across ten benchmarks in four different Re-ID settings, offering enhanced flexibility and user-friendliness. ChatReID provides a scalable, practical solution for real-world person Re-ID applications, enabling effective multi-modal interaction and fine-grained identity discrimination.
1 Introduction
Figure 1: Illustration of ChatReID’s capabilities across different person re-identification tasks. (a) shows the Standard Person Re-ID task, (b) shows Visible-infrared Person Re-ID task, (c) shows Text-to-image Re-ID task, and (d) shows Adaptive Multi-Modal Person Retrieval task.
Person re-identification (Re-ID) is a crucial task aimed at identifying individuals across different camera views using multi-modal query information. It plays a vital role in human-centric intelligent systems with broad practical applications in crowd management, smart city infrastructure, and security analytics. In recent years, large vision-language models (LVLMs) have made substantial advancements, showcasing remarkable performance by leveraging extensive world knowledge and an enhanced ability to interpret human instructions. For example, Qwen2-VL [37] introduces Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-RoPE) to improve cross-modal fusion, significantly enhancing capabilities in image and video understanding. Although some of them also have the ability to distinguish different images, they can hardly perceive fine-grained information to solve person Re-ID.
Currently, many LVLM-based approaches for person Re-ID can be broadly categorized into three main types. 1) Cross-Modal Pedestrian Features Learning: These approaches leverage the alignment capabilities of LVLMs between text and image embeddings by incorporating template language descriptions, such as “A photo of []” or “A person with [].” 2) Text-to-Image Person Re-ID: These approaches focus on retrieving target identity images using textual descriptions as queries, leveraging the alignment capabilities of LVLMs between text and image features. 3) Unified Multi-Task Framework: These methods exploit the multimodal reasoning capabilities of LVLMs to integrate different Re-ID tasks within a unified framework, facilitating a cohesive Re-ID system. However, all of them preclude the adoption of the widely recognized LVLM-based VQA inference format. It is difficult to extend these models to common application scenarios.
In this paper, we concentrate on exploring complex-content reasoning and multi-modal retrieval capabilities of LVLMs to tackle the aforementioned limitations. We develop a versatile ‘one-for-all’ person Re-ID framework, ChatReID, to interactively ask the machine to help re-identify a person given arbitrary, freeform, and necessary descriptions or clues. An illustration of some functionalities is shown in Figure 1, which demonstrates the capability to perform VQA inference format person Re-ID as well as adaptive multi-modal person retrieval. Nevertheless, achieving this goal presents significant challenges without large-scale pre-training. This complexity stems primarily from two factors: 1) Fine-Grained Identity-Level Retrieval: Unlike general image retrieval tasks, person Re-ID demands fine-grained, identity-level differentiation. While existing LVLMs can easily distinguish between broad categories (e.g., cats and dogs), they lack the specialized pre-training required to recognize individual identities based on nuanced visual or textual cues. 2) Diverse Re-ID Task Settings: Different Re-ID tasks focus on unique sets of visual features, which complicates integrating them effectively within a single model and evaluating performance using a unified VQA inference format.
To achieve this goal, we design a Hierarchical Progressive Tuning (HPT) strategy, which can be divided into three training stages: incremental pre-training for the ReID-specific vision encoder, scaling up pedestrian discrimination capability via multi-task joint training, and enhancing the pedestrian retrieval capability via post-training the decoder. Through HPT, the model’s inherent matching and reasoning capabilities are gradually enhanced to achieve fine-grained identity discrimination. In Stage 1, we perform incremental pre-training of the vision encoder to help the vision encoder understand identity-specific features of pedestrians. In Stage 2, we conduct joint training with seven tasks, enabling LVLMs to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes. In Stage 3, we design instruction fine-tuning for three specific application scenarios to enhance the practicality of ChatReID and deepen the model’s understanding of the person Re-ID task objective. As shown in Figure 2, extensive experiments demonstrate that our method achieves SOTA results across ten benchmarks in four different person Re-ID settings.
Figure 2: Performance across ten benchmarks in four different person Re-ID settings. This demonstrates ChatReID’s robustness and effectiveness across diverse Re-ID challenges.
Our paradigm has several notable characteristics:
$\bullet$ Leveraging LVLM Retrieval Capabilities. Through a Hierarchical Progressive Tuning strategy, our model fully develops the fine-grained identity discrimination capabilities of LVLMs, closely aligned with real-world application requirements.
$\bullet$ User-Friendly Design. By utilizing a VQA inference format, ChatReID makes complex person re-identification tasks accessible, enabling users to handle these tasks effectively without extensive technical expertise.
$\bullet$ Enhanced Flexibility. ChatReID can not only support traditional Re-ID tasks but also allow users to freely combine input modalities or adjust the sequence of inputs, resulting in a more flexible and accurate person Re-ID system suited to practical applications.
2 Related Work
2.1 Traditional Person Re-ID
Person re-identification (Re-ID) is a fundamental task in computer vision, which aims to match the same individual across different camera views based on visual features. Recent studies in person Re-ID carefully designed settings and developed models to tackle every specific scenario. Standard person Re-ID [59, 60, 26, 34, 14, 52], which aims to match individuals across cameras based on visual features. These methods distinguish pedestrian identities based on body posture and appearance. Cloth-changing Re-ID (CC Re-ID) [27, 4, 20, 15, 10] is a more challenging variant where individuals change their clothing between camera views. It assists the model in extracting non-clothing information for identity determination. CSSC [38] introduces a framework that effectively leverages abundant semantics within pedestrian images to extract identity features. Visible-infrared person ReID (VI-ReID) methods [7, 39, 17] extract pedestrian features under low-light environments. DDAG [49] improves performance by leveraging intra-modality and cross-modality contextual cues to enhance feature discriminability and robustness to noise. Text-to-image Re-ID [29, 11] aims to identify pedestrians based on textual descriptions. It requires the model to understand and align linguistic descriptions with visual attributes. Zhao et al. [57] proposes a novel method to model multi-modal uncertainty and semantic alignment using Gaussian distributions and a cross-modal circle loss. However, different settings within person Re-ID focus on distinct visual features, making it difficult to effectively integrate these settings into a single model. Consequently, we intend to develop a versatile ‘one-for-all’ framework to interactively ask the machine to help with the person retrieval task.
2.2 LVLM-driven Person Re-ID
Large Vision-Language Models (LVLMs) [2, 37, 48, 24] have garnered significant attention in the AI community due to their impressive generalization capabilities. Recent studies have started investigating the incorporation of LVLMs into the person Re-ID paradigm. Tan et al.[35] and Yang et al.[46] primarily focuses on the text-to-image person Re-ID task. The former uses multi-modal large language models (MLLMs) to caption images according to various templates, thereby addressing issues related to the quantity and quality of textual descriptions. The latter propose a common instruction template and use features computed by MLLMs to train person Re-ID models. Instruct-ReID [13] is the first work that unifies multiple person Re-ID settings within a single model, generating task-specific instructions and combining instruction encodings with visual encodings for re-identification training. Despite significant progress in integrating LVLMs into person Re-ID, existing methods face key limitations. Firstly, they fail to fully utilize LVLMs’ perception and instruction-following abilities. Secondly, many approaches rely on rigid, template-based textual descriptions, limiting adaptability and scalability. Lastly, while some methods unify different Re-ID settings, their flexibility remains constrained, making it difficult to apply them to common scenarios. In this paper, we present a versatile ‘one-for-all’ Re-ID framework that leverages LVLMs for interactive, freeform person Re-ID.
3 Methodology
In this paper, we propose an innovative person re-identification paradigm called ChatReID, which achieves an open-ended, interactive Re-ID application system. The subsequent sections present a comprehensive overview of the proposed methodological framework, the hierarchical progressive tuning strategy, and the data engine.
Table 1: Statistics of Datasets Used in Each Training and Inference Stage. For image-based datasets, we present the number of images and IDs. For text-to-image datasets, we show the number of images and captions.
Phase Datasets #Images #IDs/Captions
Stage 1 LUPerson-NL [8] 10,683,716 433,997
OmniReID [13] 4,973,044 328,604
Stage 2 DeepChange [43] 187,000 1,121
COCAS-Real2 [51] 21,319 101
LLCM [54] 46,767 1,064
LaST [30] 228,000 10,862
Flickr 30k [50] 31,783 158,915
UFineBench [64] 33,652 90,351
SYNTH-PEDES[63] 4,791,771 312,321
Stage 3 / Infer Market1501 [58] 32,668 $1,501$
MSMT17 [40] 126,441 4,101
CUHK03 [22] 14,096 1,467
PRCC [45] 33698 221
LTCC [27] 17,138 152
SYSU-MM01 [41] 45,863 491
RegDB [25] 8,240 412
CUHK-PEDES [21] 43,264 86,528
ICFG-PEDES [5] 54,522 54,522
RSTPReid [61] 20,505 41,010
Framework. In this section, we detail the technical architecture of ChatReID. The ChatReID framework, illustrated in Figure 3, integrates the structure of Qwen2-VL [37] due to its Naive Dynamic Resolution architecture, which effectively addresses the challenges posed by variable image resolutions in person Re-ID tasks. The hierarchical tuning process consists of three stages: Incremental Pre-training aims to help the vision encoder capture identity-specific features of pedestrians, enabling it to focus on fine-grained details relevant to person Re-ID. Both the encoder and decoder are initialized with a pre-trained LVLM to leverage existing language-vision knowledge. Multi-task Joint Training develops foundational skills in person identity matching based on images, textual descriptions, and pedestrian attributes. Only the encoder developed in Stage 1 is retained. Decoder Post-training enhances ChatReID’s understanding of Re-ID objectives and improves its instruction-following capabilities for practical applications. Here, we build upon the model from Stage 2 and freeze the encoder.
Figure 3: Overview of ChatReID’s three stage hierarchical training framework.
Training Datasets. For each training stage, we collect and organize a substantial number of person Re-ID datasets to construct the instruction tuning datasets used for training. The statistics of the datasets used in each stage are shown in Table 1. In Stage 1, we collect and organize two large-scale datasets, LUPerson-NL [8] and OmniReID [13], to construct a person image caption instruction dataset with over 5 million samples. In Stage 2, we expand our dataset collection to include a range of public datasets from various Re-ID settings. These include DeepChange [43], COCAS-Real2 [51], and LaST [30] for standard person Re-ID and CC Re-ID; LLCM [54] for VI-ReID; and UFineBench [64] and SYNTH-PEDES [63] for text-to-image Re-ID, totaling 5,340,292 images. Using these datasets, we constructed a joint training instruction dataset comprising 3 million samples across seven training tasks. In Stage 3, we gather ten widely used benchmarks across four Re-ID settings to create a comprehensive instruction dataset with 500,000 samples and three specific training objectives. These benchmarks include Market1501 [58], MSMT17 [40], and CUHK03 [22] for standard person Re-ID; PRCC [45] and LTCC [27] for CC Re-ID; SYSU-MM01 [41] and RegDB [25] for VI-ReID; and CUHK-PEDES [21], ICFG-PEDES [5], and RSTPReid [61] for text-to-image Re-ID. During the training phase, we utilized only the training subsets of these benchmarks, while the inference phase involved evaluation using their respective test sets.
3.1 Incremental Pre-training
3.1.1 Data Engine
In this stage, we utilize over 5M image-text instruction pairs. We begin by collecting and processing two large-scale person Re-ID datasets, where many images initially exhibited issues like blocking, blurring, and the presence of multiple individuals, rendering them unsuitable for pre-training. To address this, we performed post-processing to reduce noise and enhance data quality. The data construction process involved several key steps: Image Filtering by Quality. We filter images based on size and resolution to address blurriness and improve overall quality. Image Filtering by Attributes. We employ GPT-4o [1] to predict pedestrian attributes. If the model failed to recognize the pedestrian’s clothing. These are marked as occluded and removed from pre-training. Image Captioning. Using GPT-4o, we generated high-quality, diverse textual descriptions for each image. Multiple descriptions are created for each image with varying syntactic structures to enhance variability and richness.
3.1.2 The Vision Encoder pre-training
We conduct full-parameter training on both the vision encoder and language decoder of Qwen2-VL 2b. In Stage 1, full-parameter training enhances the encoder’s focus on pedestrian features, while training the decoder ensures that captions accurately reflect pedestrian attributes. The patch merger further guides the encoder’s training effectively. We standardize images with resolutions below 256x128, resizing them to 256x128 while leaving higher-resolution images unchanged. This adjustment leverages our experience with pedestrian tasks, and moderate image compression enhances training efficiency.
3.2 Multi-task Joint Training
After completing Stage 1, we obtain an encoder capable of focusing on the visual features specific to pedestrians. However, at this stage, the encoder-decoder still lacks the ability to distinguish individual identities. To address this, we implement a multi-task joint training strategy and an enhanced data engine to incorporate comprehensive pedestrian identification knowledge into ChatReID.
Figure 4: Multi-Task Design Diagram for joint training in Stage2. Concretely, we conduct seven distinct matching and retrieval tasks between text and image modalities, encouraging LVLMS to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes.
3.2.1 Multi-task Design
Multi-Task Design Diagram as shown in Figure 4. We structure the pedestrian discrimination task into three objectives: image-to-image retrieval, text-to-image retrieval, and attribute-based retrieval. Image-to-image retrieval includes standard person Re-ID, CC Re-ID, and VI-ReID. Text-to-image retrieval is used to match individuals based on textual descriptions. Attribute-Based Retrieval focuses on distinguishing individuals based on specific pedestrian attributes, which play a crucial role in person retrieval. Additionally, our multi-task design incorporates varying levels of input complexity: one-to-one pairing, one-to-many retrieval, and many-to-many retrieval. This diverse range of task complexities is essential for balancing training difficulty and achieving robust results.
Image-to-image One-to-One Matching. We sample two images from the dataset and query whether the pedestrians shown in both images belong to the same identity. To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the sampled images to belong to the same identity.
Image-to-image Many-to-Many Retrieval. We sample an image as the query image and a set of $N$ images as the gallery, where $N$ includes $n$ images that share the same identity as the query image. The task is to assess whether the model can accurately locate all images corresponding to the query identity. Both $N$ and $n$ vary dynamically.
Text-to-image One-to-One Matching. We sample an image and a text description from the dataset and query if they are a matching pair. To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the samples to belong to the same identity.
Text-to-image Many-to-One Retrieval. We sample an image and multiple text descriptions from the dataset, then query which caption accurately describes the query image.
Text-to-image One-to-Many Retrieval. We sample multiple images and a single text description from the dataset, then query which image accurately matches the description.
Attribute-level Commonality and Uniqueness Comparison. We sample two images and query the model to explain the similarities and distinctions in the identifiable attributes of the pedestrians shown. This approach enhances the model’s comprehension of pedestrian attributes. In this case, ground truth data is derived from a text-to-image dataset with attribute annotations. We manually extract shared and unique attributes from the two images to serve as supervisory information during training.
Attribute Annotations Prediction. We sample an image and query the model to predict its attribute annotations (e.g., gender, clothing color, etc.).
Instead of manual text query inputs, we create a series of multi-turn dialogue data$(X_{1}^{q},X_{1}^{a},\cdots,X_{T}^{q},X_{T}^{a})$, with T representing the complete count of dialogue exchange, addressing the divergence in multi-task interpretation between humans and large language models. For each individual task, the query format is altered, and only one instance is provided in the text.
3.2.2 Data Engine
The current LLM model exhibits superior text-to-image retrieval capabilities compared to image-to-image retrieval in the person Re-ID task. Consequently, we augment the data volume for the image-to-image tasks to significantly enhance its matching ability during this training stage. The two image-to-image tasks comprise approximately 25% to 30% of the overall data volume, whereas the remaining tasks each constitute roughly 10%.
Image-image data and text-to-image data construction. The quantity of sampled gallery images for the Many-to-Many Retrieval task adheres to a uniform distribution $N\sim\text{Uniform}(a,b)$. This primarily aims to complicate the learning process and to inhibit the model from excessively depending on selecting responses from a particular gallery size. The distribution of the response $P(n)$ is obtained based on inverse transform sampling. This approach allows us to balance the probability that each number falls within different intervals, ensuring that the probability of recording any number eventually is uniform across all numbers. The probability $P(n=k)$ is given by:
$P(n=k)=\frac{1}{k-1},\quad\text{for }k\in{2,3,\dots,N}$ (1)
This formulation ensures that each number in the range has an equal chance of being recorded, leading to a uniform coverage of all possible values.This primarily aims to prevent the model from acquiring patterns in the distribution of responses, hence diminishing the complexity of learning.
Attribute-based data construction. The SYNTH-PEDES dataset includes precise annotations of pedestrian attributes to assist with the attribute-based task. Initially, we clean the templated pedestrian attribute annotations to isolate just the properties that are present in the reported pedestrian. In the task of attribute-based generation of image comparison descriptions, we do a preliminary comparison of the sampled pedestrian attributes and utilize the results as ground truth input for the model.
3.2.3 Joint Training
We carry over the encoder from Stage 1 but replace the decoder to continue full-parameter training. The previous decoder’s strong captioning focus could hinder outcomes in this stage, so we use a freshly pre-trained decoder instead. During this stage, we adjust the size of input images to accommodate a large number of gallery images. Images larger than 384$\times$192 were rescaled to 384$\times$192, while images below this threshold were left unaltered.
3.3 Decoder Post-training
After compressing the general pedestrian discrimination information via the above two stages, ChatReID has foundational skills in person identity matching. During this stage, we fine-tune the language decoder to improve the practicality of GRP, deepening the model’s understanding of Re-ID objectives and further enhancing its instruction-following abilities for effective real-world deployment.
3.3.1 Data Engine
We rethink the assessment of the person Re-ID problem and formulate a question construction in the style of Visual Question Answering (VQA) utilizing a multi-modal large language model from a practical application perspective. The instruction tuning data amounts to 500K.
Person Re-ID. The most common type of query involves providing a text or image query alongside a gallery set, with the goal of identifying images in the gallery that match the query’s identity. However, for current LVLMs, retrieving an indefinite-length sequence of images from an indefinite gallery set based on specific criteria is highly challenging. To address this, we transform this complex retrieval task into a multi-turn best-choice problem through engineering enhancements. Specifically, at each step, the model selects the most similar image, making a locally optimal choice, which simplifies the problem and reduces computational complexity.
Attribute-Based Person Retrieval. In real-world applications, providing a fully detailed and precise query, whether text or image, is often difficult and costly. Users typically possess only fragmented textual information, such as ”a person wearing a blue top.” To accommodate this limitation, we developed a retrieval objective based on specific pedestrian attributes, enabling effective retrieval even when the query contains only partial details. Given a gallery set and a query with accurate pedestrian attribute descriptions, the model identifies the gallery image that best matches the query information.
Adaptive Multi-Modal Person Retrieval. Finally, in practical scenarios, multi-turn retrieval based on limited query information often results in a set of responses with relatively low accuracy. As retrieval continues, users typically gather more query details to improve subsequent searches. For example, beginning with attribute-based retrieval, if the correct match appears among the initial results, the user can use that match as a new query to refine further results. Alternatively, users can start with image-based retrieval and add textual details to address complexities such as clothing changes or long-term recognition tasks. By integrating multi-modal information, this approach significantly improves both the flexibility and accuracy of pedestrian re-identification, advancing its real-world applicability. To support this, we design a training objective for handling mixed queries with multi-modal inputs.
3.3.2 Post-training
We inherit the full structure from Stage 2, training with the encoder frozen. With pedestrian recognition skills already developed, this stage focuses on enhancing the model’s instruction-following capabilities. The input image resolution is preserved without any adjustments.
4 Experiments
4.1 Experimental Setups
Datasets. We adhere to the datasets used in Stage 3, using only the test sets. We utilize the commonly used metrics in person Re-ID, such as Cumulative Match Characteristic (CMC) and Mean Average Precision (mAP). Results are presented in terms of mAP/Rank-1.
Evaluation strategy. Evaluating LVLMs for person Re-ID in a VQA format presents two main challenges. First, the large gallery sets in person Re-ID datasets exceed the token limits of current LVLMs. Second, retrieving an indefinite sequence of images based on specific criteria from a vast gallery is difficult for these models. To address these issues, we implement several engineering optimizations. We begin by training a base model to filter the gallery, calculating feature similarity between the query image and gallery images. Only images with similarity above a threshold $\tau$ are retained, significantly reducing gallery size and improving evaluation efficiency without affecting accuracy. Additionally, as described in Stage 3, we convert this retrieval task into a multi-turn best-choice problem. Finally, we sample non-overlapping images from the filtered gallery and pair each selected image with the query and rank them based on response confidence, producing the final similarity list.
Table 2: Comparison of ChatReID with SOTA methods across three standard person Re-ID datasets.
Methods Market1501 MSMT17 CUHK03
mAP Rank-1 mAP Rank-1 mAP Rank-1
TransReID [12] 86.8 94.4 61.0 81.8 - -
SAN [19] 88.0 96.1 - - 76.4 80.1
HumanBench [36] 89.5 - 69.1 - 77.7 -
PASS [62] 93.0 96.8 71.8 88.2 - -
IRM [13] 93.5 96.5 72.4 86.9 85.4 86.5
ChatReID 96.4 97.2 87.5 90.1 91.3 92.7
Table 3: Comparison of ChatReID with SOTA methods across two CC Re-ID datasets.
Methods LTCC PRCC
mAP Rank-1 mAP Rank-1
HACNN [23] 26.7 60.2 - 21.8
RGA-SC [56] 27.5 65.0 - 42.3
PCB [33] 30.6 65.1 38.7 41.8
IANet [16] 31.0 63.7 45.9 46.3
CAL [9] 40.8 74.2 - -
TransReID [12] - - 44.2 -
IRM [13] 52.0 75.8 52.3 54.2
ChatReID 76.3 82.7 75.6 80.2
4.2 Experimental Results
Standard Person Re-ID. As shown in Tab. 2, ChatReID achieves impressive results across all standard person Re-ID datasets, particularly excelling on challenging datasets. The most notable improvement is observed on the complex and large-scale MSMT17 dataset, where ChatReID achieves a 15.1% increase in mAP. This substantial gain highlights the strength of our three-stage tuning strategy, which effectively extracts robust pedestrian-specific features, thereby improving model resilience in demanding scenarios. On the simpler Market1501 dataset, the performance gain is smaller, with a 2.9% increase in mAP. This modest improvement is expected, as existing methods have already achieved high accuracy on this dataset. Nonetheless, ChatReID’s measurable boost in performance demonstrates the consistent effectiveness of our approach, even when applied to datasets with strong baseline results.
CC Re-ID. The results in Tab. 3 indicate that ChatReID demonstrates significant mAP enhancements over IRM on both the LTCC and PRCC datasets, with increases of 24.3% and 23.3% in mAP, respectively. Although both IRM and ChatReID incorporate several person Re-ID settings, ChatReID outperforms for two primary reasons. First, all three stages of our tuning are based on image-text pairs, where textual information aids the visual encoder in learning variations in clothing more effectively. Second, we introduced innovative attribute-based training tasks in Stage 2, which enhance the model’s ability to recognize cloth-changing scenarios.
Table 4: Comparison of ChatReID with SOTA methods across two VI-ReID datasets.
Methods SYSU-MM01 RegDB
mAP Rank-1 mAP Rank-1
DART [44] 66.3 68.7 75.7 83.6
CAL [9] 66.9 69.6 79.1 85.0
MPANet [42] 68.2 70.6 80.7 82.8
MMN [55] 66.9 70.6 84.1 91.6
DCLNet [31] 65.3 70.8 74.3 81.2
MAUM [18] 68.8 71.7 85.1 87.9
DEEN [53] 71.8 74.7 85.1 91.1
ChatReID 83.6 86.8 95.8 96.5
Table 5: Ablation Study. Performance comparison of different stage combinations in our hierarchical progressive learning strategy across ten benchmarks.
Methods Standard ReID CC-ReID VI-ReID T2I-ReID Avg.
Market1501 MSMT17 CUHK03 PRCC LTCC SYSU-MM01 RegDB CUHK-PEDES. ICFG-PEDES RSTPReid
Stage 3 59.2 41.5 60.3 31.2 35.4 11.4 12.7 65.1 49.9 43.5 41.0
Stage 2 + Stage 3 92.1 79.5 85.7 60.2 55.1 76.2 74.7 77.8 62.5 71.4 73.5
ChatReID 96.4 87.5 91.3 76.3 75.6 83.6 95.8 80.1 70.5 73.1 83.0
Table 6: Comparison of ChatReID with SOTA methods across three T2I Re-ID datasets.
Methods CUHK-PEDES ICFG-PEDES RSTPReid
mAP Rank-1 mAP Rank-1 mAP Rank-1
IRRA [18] 66.1 73.4 38.1 63.5 - 60.2
RDE [28] 67.6 75.9 40.1 67.7 50.9 65.4
WoRA [32] 67.2 76.4 42.6 87.5 52.5 66.9
RaSa [3] 69.4 76.5 41.3 65.3 52.3 67.0
APTM [47] 66.9 76.5 41.2 68.5 - 67.5
MARS [6] 71.7 77.6 44.9 67.6 52.9 67.6
IRM [13] 66.5 74.2 - - - -
ChatReID 80.1 83.8 70.5 72.9 73.1 75.0
VI-ReID. As shown in Tab. 4, ChatReID also demonstrates significant improvements on VI-ReID datasets. Specifically, it achieves an 11.8% mAP increase on SYSU-MM01 and a 10.7% mAP increase on RegDB. These gains are primarily due to our approach in Stage 2, where we treated the VI-ReID task as an image-image matching task and slightly increased the proportion of VI data in the tuning process. This adjustment allowed the model to learn VI matching more effectively, resulting in enhanced performance in cross-modality Re-ID tasks.
Text-to-Image Person Re-ID. As shown in Tab. 6, ChatReID achieves remarkable gains in text-to-image person Re-ID, especially on the ICFG-PEDES and RSTPReid datasets, where prior performance was relatively low. ChatReID reaches an mAP of 70.5% on ICFG-PEDES, marking a 25.6% improvement, and an mAP of 73.1% on RSTPReid, a 20.2% increase. These significant gains can be attributed to two key factors. First, existing LVLMs already possess a baseline level of text-to-image matching capability, as evidenced in our ablation studies. Second, throughout all three training stages, we utilize a dataset structure based on image-text pairs, which substantially enhances the model’s ability to learn pedestrian-specific text-to-image matching.
Attribute-Based Person Retrieval. To our knowledge, there is currently no straightforward and effective method for the Attribute-Based Person Retrieval task. Existing text-to-image person Re-ID methods primarily focus on retrieving images based on complete textual descriptions rather than specific attributes. To address this gap, we introduced attribute-based retrieval as a training objective in Stage 3 of our model. To assess the effectiveness of our approach, we reorganized the dataset by randomly sampling $N$ images from an attribute-annotated dataset. Our goal was to retrieve images that matched specified attributes, such as clothing or accessories. Experimental results indicate that ChatReID can efficiently retrieve images based on attribute criteria, underscoring its practical value for real-world applications. Additional experimental results are presented in the supplementary material.
Adaptive Multi-Modal Person Retrieval. Considering practical application needs, we design a training task for Adaptive Multi-Modal Person Retrieval. To evaluate the model’s performance on this task, we conducted experiments that included additional textual information as auxiliary input for image-image retrieval tasks. The experimental results demonstrate that ChatReID can efficiently handle person retrieval with multi-modal inputs and that incorporating multi-modal information enhances performance in person Re-ID task. Detailed results and analyses are provided in the supplementary material.
4.3 Ablation Study
To validate the effectiveness of the three-stage hierarchical progressive tuning strategy, we compare experimental results from different stage combinations. The results are shown in Tab. 5. We first report results from the application of Stage 3 training directly to the pre-trained Qwen2-VL model. The results show that while the pre-trained Qwen2-VL achieves some baseline performance, the results are suboptimal, indicating that current LVLMs possess limited inherent image-image and image-text pedestrian matching capabilities but require specialized training for optimal performance. Performance on the VI-ReID task was particularly low, likely due to the limited proportion of low-illumination data in the training set, which is insufficient for training a robust model. In our experiments, we found that combining only Stage 1 and Stage 3 made it difficult for the model to converge; thus, we did not report those results. However, adding Stage 2 significantly improved model performance, achieving an average mAP of 73.5%, which demonstrates practical utility. This validates the effectiveness of our Stage 2 training design. When Stage 1 training is also included, the model’s performance further improves, reaching SOTA results across ten benchmarks. This final enhancement reinforces the overall effectiveness of our hierarchical three-stage tuning strategy.
5 Conclusion
This paper introduces a novel framework for person re-identification (Re-ID), ChatReID, that leverages the advanced capabilities of large vision-language models (LVLMs). Unlike traditional methods, which are limited by rigid templates and lack of adaptability, our system enables freeform, user-driven queries, offering greater flexibility and natural interaction. By implementing a Hierarchical Progressive Tuning (HPT) strategy, we progressively enhance the model’s ability to achieve fine-grained identity-level retrieval. Our approach, combining Incremental Pre-training, Multi-task Joint Training, and Decoder Post-training , delivers SOTA performance across various Re-ID benchmarks. The system’s VQA inference format simplifies complex Re-ID tasks, making it accessible to non-experts, while its flexibility allows dynamic input combinations and adjustments. Our work establishes a strong foundation for future advancements in person Re-ID, showcasing the potential of LVLMs in practical applications.

### v2

@@
 Abstract
-Person re-identification (Re-ID) is a critical task in human-centric intelligent systems, enabling consistent identification of individuals across different camera views using multi-modal query information.
-Recent studies have successfully integrated LVLMs with person Re-ID, yielding promising results.
-However, existing LVLM-based methods face several limitations.
-They rely on extracting textual embeddings from fixed templates, which are used either as intermediate features for image representation or for prompt tuning in domain-specific tasks.
-Furthermore, they are unable to adopt the VQA inference format, significantly restricting their broader applicability.
-In this paper, we propose a novel, versatile, one-for-all person Re-ID framework, ChatReID.
-Our approach introduces a Hierarchical Progressive Tuning (HPT) strategy, which ensures fine-grained identity-level retrieval by progressively refining the model’s ability to distinguish pedestrian identities.
-Extensive experiments demonstrate that our approach outperforms SOTA methods across ten benchmarks in four different Re-ID settings, offering enhanced flexibility and user-friendliness.
-ChatReID provides a scalable, practical solution for real-world person Re-ID applications, enabling effective multi-modal interaction and fine-grained identity discrimination.
+Person re-identification (Re-ID) is a crucial task in computer vision, aiming to recognize individuals across non-overlapping camera views.
+While recent advanced vision-language models (VLMs) excel in logical reasoning and multi-task generalization, their applications in Re-ID tasks remain limited.
+They either struggle to perform accurate matching based on identity-relevant features or assist image-dominated branches as auxiliary semantics.
+In this paper, we propose a novel framework ChatReID, that shifts the focus towards a text-side-dominated retrieval paradigm, enabling flexible and interactive re-identification.
+To integrate the reasoning abilities of language models into Re-ID pipelines, We first present a large-scale instruction dataset, which contains more than 8 million prompts to promote the model fine-tuning.
+Next.
+we introduce a hierarchical progressive tuning strategy, which endows Re-ID ability through three stages of tuning, i.e., from person attribute understanding to fine-grained image retrieval and to multi-modal task reasoning.
+Extensive experiments across ten popular benchmarks demonstrate that ChatReID outperforms existing methods, achieving state-of-the-art performance in all Re-ID tasks.
+More experiments demonstrate that ChatReID not only has the ability to recognize fine-grained details but also to integrate them into a coherent reasoning process.
 1 Introduction
-Figure 1: Illustration of ChatReID’s capabilities across different person re-identification tasks.
-(a) shows the Standard Person Re-ID task, (b) shows Visible-infrared Person Re-ID task, (c) shows Text-to-image Re-ID task, and (d) shows Adaptive Multi-Modal Person Retrieval task.
-Person re-identification (Re-ID) is a crucial task aimed at identifying individuals across different camera views using multi-modal query information.
-It plays a vital role in human-centric intelligent systems with broad practical applications in crowd management, smart city infrastructure, and security analytics.
-In recent years, large vision-language models (LVLMs) have made substantial advancements, showcasing remarkable performance by leveraging extensive world knowledge and an enhanced ability to interpret human instructions.
-For example, Qwen2-VL [37] introduces Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-RoPE) to improve cross-modal fusion, significantly enhancing capabilities in image and video understanding.
-Although some of them also have the ability to distinguish different images, they can hardly perceive fine-grained information to solve person Re-ID.
-Currently, many LVLM-based approaches for person Re-ID can be broadly categorized into three main types.
-1) Cross-Modal Pedestrian Features Learning: These approaches leverage the alignment capabilities of LVLMs between text and image embeddings by incorporating template language descriptions, such as “A photo of []” or “A person with [].” 2) Text-to-Image Person Re-ID: These approaches focus on retrieving target identity images using textual descriptions as queries, leveraging the alignment capabilities of LVLMs between text and image features.
-3) Unified Multi-Task Framework: These methods exploit the multimodal reasoning capabilities of LVLMs to integrate different Re-ID tasks within a unified framework, facilitating a cohesive Re-ID system.
-However, all of them preclude the adoption of the widely recognized LVLM-based VQA inference format.
-It is difficult to extend these models to common application scenarios.
-In this paper, we concentrate on exploring complex-content reasoning and multi-modal retrieval capabilities of LVLMs to tackle the aforementioned limitations.
-We develop a versatile ‘one-for-all’ person Re-ID framework, ChatReID, to interactively ask the machine to help re-identify a person given arbitrary, freeform, and necessary descriptions or clues.
-An illustration of some functionalities is shown in Figure 1, which demonstrates the capability to perform VQA inference format person Re-ID as well as adaptive multi-modal person retrieval.
-Nevertheless, achieving this goal presents significant challenges without large-scale pre-training.
-This complexity stems primarily from two factors: 1) Fine-Grained Identity-Level Retrieval: Unlike general image retrieval tasks, person Re-ID demands fine-grained, identity-level differentiation.
-While existing LVLMs can easily distinguish between broad categories (e.g., cats and dogs), they lack the specialized pre-training required to recognize individual identities based on nuanced visual or textual cues.
-2) Diverse Re-ID Task Settings: Different Re-ID tasks focus on unique sets of visual features, which complicates integrating them effectively within a single model and evaluating performance using a unified VQA inference format.
-To achieve this goal, we design a Hierarchical Progressive Tuning (HPT) strategy, which can be divided into three training stages: incremental pre-training for the ReID-specific vision encoder, scaling up pedestrian discrimination capability via multi-task joint training, and enhancing the pedestrian retrieval capability via post-training the decoder.
-Through HPT, the model’s inherent matching and reasoning capabilities are gradually enhanced to achieve fine-grained identity discrimination.
-In Stage 1, we perform incremental pre-training of the vision encoder to help the vision encoder understand identity-specific features of pedestrians.
-In Stage 2, we conduct joint training with seven tasks, enabling LVLMs to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes.
-In Stage 3, we design instruction fine-tuning for three specific application scenarios to enhance the practicality of ChatReID and deepen the model’s understanding of the person Re-ID task objective.
-As shown in Figure 2, extensive experiments demonstrate that our method achieves SOTA results across ten benchmarks in four different person Re-ID settings.
-Figure 2: Performance across ten benchmarks in four different person Re-ID settings. This demonstrates ChatReID’s robustness and effectiveness across diverse Re-ID challenges.
-Our paradigm has several notable characteristics:
-$\bullet$ Leveraging LVLM Retrieval Capabilities.
-Through a Hierarchical Progressive Tuning strategy, our model fully develops the fine-grained identity discrimination capabilities of LVLMs, closely aligned with real-world application requirements.
-$\bullet$ User-Friendly Design.
-By utilizing a VQA inference format, ChatReID makes complex person re-identification tasks accessible, enabling users to handle these tasks effectively without extensive technical expertise.
-$\bullet$ Enhanced Flexibility.
-ChatReID can not only support traditional Re-ID tasks but also allow users to freely combine input modalities or adjust the sequence of inputs, resulting in a more flexible and accurate person Re-ID system suited to practical applications.
+Figure 1: Performance comparisons across ten benchmarks on four person Re-ID tasks show that our ChatReID significantly outperforms previous state-of-the-art methods, demonstrating its superior robustness and effectiveness.
+Figure 2: (a) and (b) show the results of standard and cloth-changing(CC) Re-ID using GPT-4o and Qwen2-VL without tuning.
+(c)–(g) illustrate ChatReID’s capabilities across various person Re-ID tasks.
+Red dialogue boxes indicate incorrect responses, while green dialogue boxes indicate correct responses.
+Best viewed in color and zoom in.
+Person re-identification (Re-ID) is a fundamental task in computer vision that aims to recognize individuals across non-overlapping camera views [55, 49, 14].
+It plays a crucial role in intelligent surveillance systems, facilitating applications in suspect tracking, crowd management, and access control.
+In the past decade, person Re-ID has achieved remarkable progress with the development of deep learning techniques [7, 50, 4], significantly relieving challenges of pose variants, illumination and occlusions.
+Recently, the emergence of vision-language models (VLMs) [1, 39, 46, 26] present new Re-ID paradigms by integrating textual descriptions with visual representations.
+They leverage the ability of image-text alignment to enrich feature representations, further promoting the performance and application range of person Re-ID.
+Depending on the usage of text information, VLM-based methods can be broadly grouped into three categories.
+(1) Auxiliary supervision using text [23, 47]: they employ text embedding to additionally supervise the feature learning or to integrate with image features.
+(2) Text-based person retrieval [37, 45]: they adopt textual descriptions as queries to retrieve images with the matched identity.
+(3) Text prompts for unified Re-ID [12, 13]: they use text prompts or language instructions to manipulate a unified model handling different Re-ID tasks.
+Despite the impressive effect, these methods still follow the traditional image-side dominated Re-ID paradigm, i.e., extracting features with feature extractors (or encoders) and then sorting them by calculating feature similarity.
+Essentially, person Re-ID is a complex reasoning process that requires analyzing a person’s appearance and biological information from an image, then repeatedly comparing different and similar points in a pair of images to determine whether their identities match.
+However, existing VLMs, such as Qwen2-VL [39] and GPT-4o [18], posses strong vision-language reasoning abilities in programming and communication, yet their potential in person Re-ID tasks remains largely unstudied.
+This raises a natural question: can we leverage the advanced reasoning capabilities of VLMs to perform text-side dominated person Re-ID?
+Intuitively, a straightforward practice is to retain the text decoder of VLMs in the pipeline and ask it to provide the index of the matched images or the similarity score between the given query and gallery images.
+However, as depicted in Fig.
+2(a) and (b), GPT-4o focuses only on appearance features, leading to incorrect matches, while Qwen2-VL produces almost illogical results.
+We argue that successful person Re-ID requires not only recognizing fine-grained details but also integrating them into a coherent reasoning process.
+Concretely, (1) person identity is inherently an intra-class semantics, meaning that person Re-ID relies on fine-grained biological or identity-related features.
+While existing VLMs can easily distinguish between broad categories (i.e., cats v.s.
+dogs) or major attributes (i.e., clothing colors), they lack the specialized tuning needed to identify individuals based on subtle visual or textual clues.
+(2) In addition, person Re-ID is an open-set task, where the correct matches can vary depending on the given query or task context.
+Unlike approaches that use text prompts as conditions, the text-side dominated Re-ID framework requires inferring the desirable similarities through a deep and comprehensive understanding of both textual descriptions and visual content.
+In this paper, we explore complex-content reasoning and multi-modal retrieval capabilities of VLMs and introduce ChatReID, a new perspective of using the text decoder for person Re-ID.
+Our ChatReID is a versatile ‘one-for-all’ framework to interactively ask the model to assist in re-identifying a person given arbitrary, freeform, and necessary descriptions or clues, as shown in Fig.
+2(c)-(g).
+To overcome the aforementioned challenges, we first propose a hierarchical progressive tuning (HPT) strategy, which consists of three stages.
+The first stage helps the model understand the attributes or semantics of a person.
+In the second stage, the model is guided to learn the ability of image-to-image, image-to-attribute, and text-to-image fine-grained retrieval.
+The third stage deepens the logical reasoning ability of the model between application scenario descriptions and inputs.
+Secondly, we create a large-scale instruction dataset with more than 8 million prompts, to promote the tuning of three progressive stages.
+Contributions. We summarize contributions as follows:
+  • We propose ChatReID, a novel framework for person Re-ID that introduces a text-side-dominated retrieval paradigm.
+ChatReID enables a flexible and interactive retrieval process, enhancing stronger generalization ability.
+  • We present a large-scale instruction dataset and a hierarchical progressive tuning strategy, which endows Re-ID ability through three stages of tuning, i.e., from person attribute understanding to fine-grained image retrieval and to multi-modal task reasoning.
+  • Extensive experiments on ten widely used benchmarks across four different person Re-ID tasks to evaluate the effectiveness of our model.
+ChatReID achieves state-of-the-art performance in all experiments, outperforming existing methods by a significant margin.
 2 Related Work
@@
 It assists the model in extracting non-clothing information for identity determination.
-CSSC [38] introduces a framework that effectively leverages abundant semantics within pedestrian images to extract identity features.
+CSSC [40] introduces a framework that leverages abundant semantics within pedestrian images to extract identity features.
 Visible-infrared person ReID (VI-ReID) methods [6, 41, 17] extract pedestrian features under low-light environments.
@@
 Consequently, we intend to develop a versatile ‘one-for-all’ framework to interactively ask the machine to help with the person retrieval task.
-2.2 LVLM-driven Person Re-ID
-Large Vision-Language Models (LVLMs) [2, 37, 48, 24] have garnered significant attention in the AI community due to their impressive generalization capabilities.
-Recent studies have started investigating the incorporation of LVLMs into the person Re-ID paradigm.
+2.2 VLM-driven Person Re-ID
+Vision-language models (VLMs) [1, 39, 46, 26] have garnered significant attention in the AI community due to their impressive generalization capabilities.
+Recent studies have started investigating the incorporation of VLMs into the person Re-ID paradigm.
 Tan et al.[37] and Yang et al.[44] primarily focuses on the text-to-image person Re-ID task.
 The former uses multi-modal large language models (MLLMs) to caption images according to various templates, thereby addressing issues related to the quantity and quality of textual descriptions.
-The latter propose a common instruction template and use features computed by MLLMs to train person Re-ID models.
-Instruct-ReID [13] is the first work that unifies multiple person Re-ID settings within a single model, generating task-specific instructions and combining instruction encodings with visual encodings for re-identification training.
-Despite significant progress in integrating LVLMs into person Re-ID, existing methods face key limitations.
-Firstly, they fail to fully utilize LVLMs’ perception and instruction-following abilities.
+The latter proposes a common instruction template and uses features computed by MLLMs to train person Re-ID models.
+Instruct-ReID [12] is the first work that unifies multiple person Re-ID settings within a single model, generating task-specific instructions and combining instruction encodings with visual encodings for Re-ID training.
+Despite significant progress in integrating VLMs into person Re-ID, existing methods face key limitations.
+Firstly, they fail to fully utilize VLMs’ perception and instruction-following abilities.
 Secondly, many approaches rely on rigid, template-based textual descriptions, limiting adaptability and scalability.
 Lastly, while some methods unify different Re-ID settings, their flexibility remains constrained, making it difficult to apply them to common scenarios.
-In this paper, we present a versatile ‘one-for-all’ Re-ID framework that leverages LVLMs for interactive, freeform person Re-ID.
+In this paper, we present a versatile ‘one-for-all’ Re-ID framework that leverages VLMs for interactive, freeform person Re-ID.
+Figure 3: Overview of the ChatReID framework. (a) shows the schematic of ChatReID. (b) shows the three-stage HPT strategy.
 3 Methodology
-In this paper, we propose an innovative person re-identification paradigm called ChatReID, which achieves an open-ended, interactive Re-ID application system.
-The subsequent sections present a comprehensive overview of the proposed methodological framework, the hierarchical progressive tuning strategy, and the data engine.
-Table 1: Statistics of Datasets Used in Each Training and Inference Stage.
-For image-based datasets, we present the number of images and IDs.
-For text-to-image datasets, we show the number of images and captions.
-Phase Datasets #Images #IDs/Captions
-Stage 1 LUPerson-NL [8] 10,683,716 433,997
-OmniReID [13] 4,973,044 328,604
-Stage 2 DeepChange [43] 187,000 1,121
-COCAS-Real2 [51] 21,319 101
-LLCM [54] 46,767 1,064
-LaST [30] 228,000 10,862
-Flickr 30k [50] 31,783 158,915
-UFineBench [64] 33,652 90,351
-SYNTH-PEDES[63] 4,791,771 312,321
-Stage 3 / Infer Market1501 [58] 32,668 $1,501$
-MSMT17 [40] 126,441 4,101
-CUHK03 [22] 14,096 1,467
-PRCC [45] 33698 221
-LTCC [27] 17,138 152
-SYSU-MM01 [41] 45,863 491
-RegDB [25] 8,240 412
-CUHK-PEDES [21] 43,264 86,528
-ICFG-PEDES [5] 54,522 54,522
-RSTPReid [61] 20,505 41,010
-Framework.
-In this section, we detail the technical architecture of ChatReID.
-The ChatReID framework, illustrated in Figure 3, integrates the structure of Qwen2-VL [37] due to its Naive Dynamic Resolution architecture, which effectively addresses the challenges posed by variable image resolutions in person Re-ID tasks.
-The hierarchical tuning process consists of three stages: Incremental Pre-training aims to help the vision encoder capture identity-specific features of pedestrians, enabling it to focus on fine-grained details relevant to person Re-ID.
-Both the encoder and decoder are initialized with a pre-trained LVLM to leverage existing language-vision knowledge.
-Multi-task Joint Training develops foundational skills in person identity matching based on images, textual descriptions, and pedestrian attributes.
-Only the encoder developed in Stage 1 is retained.
-Decoder Post-training enhances ChatReID’s understanding of Re-ID objectives and improves its instruction-following capabilities for practical applications.
-Here, we build upon the model from Stage 2 and freeze the encoder.
-Figure 3: Overview of ChatReID’s three stage hierarchical training framework.
-Training Datasets.
+Our ChatReID is a text-side-dominated framework for person Re-ID.
+Different from traditional methods that calculates image feature distance as similarity, ChatReID interprets the task requirements from the input text description and performs similarity inference on the given person images accordingly.
+Finally, it outputs the matched person image in textual form.
+Figure 3(a) illustrates the schematic of our ChatReID, which is primarily composed of a ViT encoder and a LLM decoder.
+In this section, we start by introducing how to fine-tune the encoder and decoder through our proposed hierarchical progressive tuning strategy (in 3.1).
+Next, we elaborate a large-scale customized instruction dataset built for fine-tuning (see 3.2).
+Lastly, we discuss the architecture details of our framework, along with the training and inference processes in Sec.
+3.3.
+3.1 Hierarchical Progressive Tuning
+Person Re-ID is a fine-grained retrieval task where identity represents highly abstract semantic information, making it difficult to achieve accurate matches using existing VLMs directly, such as Qwen2-VL [39] and GPT-4o [18] shown in Fig.
+2(a)-(b).
+However, VLMs excel in logical reasoning and multi-task generalization, offering great potential for Re-ID.
+To bridge this gap, as shown in Fig.
+3(b), we introduce a hierarchical progressive fine-tuning strategy, which gradually enhances the model’s Re-ID capability through three stages: (1) person attribute understanding, (2) fine-grained image retrieval, and (3) multi-modal task reasoning.
+3.1.1 Stage One: Person Attribtue Understanding
+Considering the diversity of person images, the first stage focuses on improving the model’s ability to understand pedestrians by learning and extracting fine-grained attributes such as gender, clothing, and posture.
+$\bullet$ Image captioning.
+Following previous studies [25, 22], we use image captioning as a foundational fine-tuning task.
+Specifically, the model is trained to generate detailed attribute descriptions based on a given person image and corresponding prompts.
+To further guide the model, we empirically incorporate several predefined key attributes into the prompt.
+This design not only encourages the model to capture more precise and informative attribute descriptions but also potentially strengthens its ability to differentiate identities.
+The prompt used is:
+“In the {person image} provided, can you give a detailed description of the pedestrian, including their gender, age range, hair, type and color of clothing, footwear type and color, posture or gait, any patterns or accessories, and any distinguishing features?”
+3.1.2 Stage Two: Fine-grained Image Retrieval
+After completing the first stage of tuning, we obtain a vision encoder that can effectively capture fine-grained person attributes.
+To further promote its ability for person re-identification, we introduce several fine-tuning tasks in the second stage to guide the model in matching identity-related features with fine-grained information.
+Specifically, we structure tasks into three categories of retrieval tasks: image-to-image, text-to-text, and image-to-attribute.
+Additionally, each type of retrieval task is designed with varying levels of complexity, including one-to-one, one-to-many, and many-to-many scenarios.
+This diverse range of task complexities is crucial for balancing training difficulty and improving the model’s robustness (see our discussions in Sec.
+4.3).
+Figure 4 illustrates seven fine-tuning tasks, including examples of prompts and responses, involved in the second stage.
+Figure 4: Multi-task Design Diagram for joint training in the stage 2.
+Concretely, we conduct seven distinct matching and retrieval tasks between text and image modalities, encouraging VLMS to acquire an initial capability for fine-grained image retrieval based on images, textual descriptions, and pedestrian attributes.
+$\bullet$ Image-to-image Retrieval.
+This task has two levels of one-to-one and one-to-many.
+For the one-to-one level, we sample two images and ask the model to determine whether the people in these two images are the same person.
+To balance the training, we set a probability of 0.5 to sample images with positive answers.
+For the one-to-many level, we randomly select one image as the query and choose a set of $N$ images, where $N$ images include $n$ images that have the same identity.
+The model is required to identify all matched images.
+During training, values of both $N$ and $n$ vary to make the task more challenging and diverse.
+$\bullet$ Text-to-image Retrieval.
+This task aims to identify the matched person image based on a text description, which has three levels of one-to-one, many-to-one, and one-to-many.
+Similar to the image-to-image retrieval task, the one-to-one level requires the model to determine if the image and the text description match or not (‘yes’ or ‘no’).
+And for the last two levels, both need to find the correctest answer from a set of options.
+$\bullet$ Image-to-attribute Retrieval.
+This task is proposed to learn to distinguish individuals based on specific attributes, and we design two specific levels.
+For the first one, two images are sampled and the model is tasked with explaining the similarities and distinctions in attributes of two pedestrians.
+It is expected to further enhance the model’s understanding ability of person attributes.
+For the second, an image is randomly selected and fed into the model to generate its attribute descriptions or annotations (e.g., gender, clothing color, etc.).
+It is similar to the first stage, with the intention to deepen and consolidate this capability of the model in the current stage.
+3.1.3 Stage Three: Multi-modal Task Reasoning
+Stage Three enhances ChatReID’s rasoning capability of Re-ID objectives and improves its instruction-following capabilities for practical applications.
+After compressing the fine-grained pedestrian discrimination information via the above two stages, ChatReID has foundational skills in person identity matching.
+In this stage, we directly adopt the training objective of 4 person Re-ID tasks.
+$\bullet$ Standard person Re-ID.
+Standard person Re-ID is a image-to-image retrieval task, where a query image and a set of gallery images are provided, and the goal is to identify the pedestrian in the gallery that matches the identity of the query, the prompt used is:
+“Given the {person image} provided, identify which of the following images show the same pedestrian as in the first image.”
+$\bullet$ CC person Re-ID. Based on the cloth-changing person Re-ID setting, the prompt used is:
+“Given the {person image} provided, select which of the {gallery} shows the same person, who may be wearing different clothing.”
+$\bullet$ VI person Re-ID. Based on the VI person Re-ID setting, the prompt used is:
+“Using the given {person image}, choose the corresponding {gallery} that matches the same pedestrian, considering both infrared and visible light images for re-identification.”
+$\bullet$ T2I person Re-ID.
+T2I person Re-ID is a text-to-image retrieval task, where a pedestrian image description is given, and the goal is to find the matching pedestrian image in the gallery based on the description, the prompt used is:
+“Select the {person image} that best matches the following description of the pedestrian’s appearance: {text} ’̇’
+To improve robustness and prevent bias from fixed-length galleries, we introduce variability in the gallery length, randomly adjusting it while ensuring at least one image matches the query identity.
+3.2 Data Engine
 For each training stage, we collect and organize a substantial number of person Re-ID datasets to construct the instruction tuning datasets used for training.
-The statistics of the datasets used in each stage are shown in Table 1.
-In Stage 1, we collect and organize two large-scale datasets, LUPerson-NL [8] and OmniReID [13], to construct a person image caption instruction dataset with over 5 million samples.
-In Stage 2, we expand our dataset collection to include a range of public datasets from various Re-ID settings.
-These include DeepChange [43], COCAS-Real2 [51], and LaST [30] for standard person Re-ID and CC Re-ID; LLCM [54] for VI-ReID; and UFineBench [64] and SYNTH-PEDES [63] for text-to-image Re-ID, totaling 5,340,292 images.
-Using these datasets, we constructed a joint training instruction dataset comprising 3 million samples across seven training tasks.
-In Stage 3, we gather ten widely used benchmarks across four Re-ID settings to create a comprehensive instruction dataset with 500,000 samples and three specific training objectives.
-These benchmarks include Market1501 [58], MSMT17 [40], and CUHK03 [22] for standard person Re-ID; PRCC [45] and LTCC [27] for CC Re-ID; SYSU-MM01 [41] and RegDB [25] for VI-ReID; and CUHK-PEDES [21], ICFG-PEDES [5], and RSTPReid [61] for text-to-image Re-ID.
-During the training phase, we utilized only the training subsets of these benchmarks, while the inference phase involved evaluation using their respective test sets.
-3.1 Incremental Pre-training
-3.1.1 Data Engine
-In this stage, we utilize over 5M image-text instruction pairs.
-We begin by collecting and processing two large-scale person Re-ID datasets, where many images initially exhibited issues like blocking, blurring, and the presence of multiple individuals, rendering them unsuitable for pre-training.
-To address this, we performed post-processing to reduce noise and enhance data quality.
+We utilize a total of 19 open-source person Re-ID datasets.
+The complete list of datasets and corresponding statistical information can be found in the supplementary materials.
+In Stage 1, we utilize over 5M image-text instruction pairs.
+To address the low-quality problem, we performed post-processing to reduce noise and enhance data quality.
 The data construction process involved several key steps: Image Filtering by Quality.
@@
 Image Captioning.
-Using GPT-4o, we generated high-quality, diverse textual descriptions for each image.
-Multiple descriptions are created for each image with varying syntactic structures to enhance variability and richness.
-3.1.2 The Vision Encoder pre-training
-We conduct full-parameter training on both the vision encoder and language decoder of Qwen2-VL 2b.
-In Stage 1, full-parameter training enhances the encoder’s focus on pedestrian features, while training the decoder ensures that captions accurately reflect pedestrian attributes.
-The patch merger further guides the encoder’s training effectively.
-We standardize images with resolutions below 256x128, resizing them to 256x128 while leaving higher-resolution images unchanged.
-This adjustment leverages our experience with pedestrian tasks, and moderate image compression enhances training efficiency.
-3.2 Multi-task Joint Training
-After completing Stage 1, we obtain an encoder capable of focusing on the visual features specific to pedestrians.
-However, at this stage, the encoder-decoder still lacks the ability to distinguish individual identities.
-To address this, we implement a multi-task joint training strategy and an enhanced data engine to incorporate comprehensive pedestrian identification knowledge into ChatReID.
-Figure 4: Multi-Task Design Diagram for joint training in Stage2.
-Concretely, we conduct seven distinct matching and retrieval tasks between text and image modalities, encouraging LVLMS to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes.
-3.2.1 Multi-task Design
-Multi-Task Design Diagram as shown in Figure 4.
-We structure the pedestrian discrimination task into three objectives: image-to-image retrieval, text-to-image retrieval, and attribute-based retrieval.
-Image-to-image retrieval includes standard person Re-ID, CC Re-ID, and VI-ReID.
-Text-to-image retrieval is used to match individuals based on textual descriptions.
-Attribute-Based Retrieval focuses on distinguishing individuals based on specific pedestrian attributes, which play a crucial role in person retrieval.
-Additionally, our multi-task design incorporates varying levels of input complexity: one-to-one pairing, one-to-many retrieval, and many-to-many retrieval.
-This diverse range of task complexities is essential for balancing training difficulty and achieving robust results.
-Image-to-image One-to-One Matching.
-We sample two images from the dataset and query whether the pedestrians shown in both images belong to the same identity.
-To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the sampled images to belong to the same identity.
-Image-to-image Many-to-Many Retrieval.
-We sample an image as the query image and a set of $N$ images as the gallery, where $N$ includes $n$ images that share the same identity as the query image.
-The task is to assess whether the model can accurately locate all images corresponding to the query identity.
-Both $N$ and $n$ vary dynamically.
-Text-to-image One-to-One Matching.
-We sample an image and a text description from the dataset and query if they are a matching pair.
-To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the samples to belong to the same identity.
-Text-to-image Many-to-One Retrieval. We sample an image and multiple text descriptions from the dataset, then query which caption accurately describes the query image.
-Text-to-image One-to-Many Retrieval. We sample multiple images and a single text description from the dataset, then query which image accurately matches the description.
-Attribute-level Commonality and Uniqueness Comparison.
-We sample two images and query the model to explain the similarities and distinctions in the identifiable attributes of the pedestrians shown.
-This approach enhances the model’s comprehension of pedestrian attributes.
-In this case, ground truth data is derived from a text-to-image dataset with attribute annotations.
-We manually extract shared and unique attributes from the two images to serve as supervisory information during training.
-Attribute Annotations Prediction. We sample an image and query the model to predict its attribute annotations (e.g., gender, clothing color, etc.).
-Instead of manual text query inputs, we create a series of multi-turn dialogue data$(X_{1}^{q},X_{1}^{a},\cdots,X_{T}^{q},X_{T}^{a})$, with T representing the complete count of dialogue exchange, addressing the divergence in multi-task interpretation between humans and large language models.
-For each individual task, the query format is altered, and only one instance is provided in the text.
-3.2.2 Data Engine
-The current LLM model exhibits superior text-to-image retrieval capabilities compared to image-to-image retrieval in the person Re-ID task.
-Consequently, we augment the data volume for the image-to-image tasks to significantly enhance its matching ability during this training stage.
-The two image-to-image tasks comprise approximately 25% to 30% of the overall data volume, whereas the remaining tasks each constitute roughly 10%.
-Image-image data and text-to-image data construction.
-The quantity of sampled gallery images for the Many-to-Many Retrieval task adheres to a uniform distribution $N\sim\text{Uniform}(a,b)$.
-This primarily aims to complicate the learning process and to inhibit the model from excessively depending on selecting responses from a particular gallery size.
+Using GPT-4o [18], we generated high-quality, diverse textual descriptions for each image.
+In Stage 2, we use seven datasets from various Re-ID settings.
+We constructed a joint training instruction dataset comprising 3M samples.
+Given that the current LLM model demonstrates superior text-to-image retrieval capabilities compared to image-to-image retrieval in the person Re-ID task, we augment the data volume for the image-to-image tasks.
+The two image-to-image tasks account for approximately 25% to 30% of the overall data volume, while the remaining tasks each contribute roughly 10%.
+Gallery image sampling The quantity of sampled gallery images for the Many-to-Many Retrieval task adheres to a uniform distribution $N\sim\text{Uniform}(a,b)$.
 The distribution of the response $P(n)$ is obtained based on inverse transform sampling.
@@
 $P(n=k)=\frac{1}{k-1},\quad\text{for }k\in{2,3,\dots,N}$ (1)
-This formulation ensures that each number in the range has an equal chance of being recorded, leading to a uniform coverage of all possible values.This primarily aims to prevent the model from acquiring patterns in the distribution of responses, hence diminishing the complexity of learning.
-Attribute-based data construction.
-The SYNTH-PEDES dataset includes precise annotations of pedestrian attributes to assist with the attribute-based task.
-Initially, we clean the templated pedestrian attribute annotations to isolate just the properties that are present in the reported pedestrian.
-In the task of attribute-based generation of image comparison descriptions, we do a preliminary comparison of the sampled pedestrian attributes and utilize the results as ground truth input for the model.
-3.2.3 Joint Training
-We carry over the encoder from Stage 1 but replace the decoder to continue full-parameter training.
+In Stage 3, we gather ten widely used benchmarks across four Re-ID settings to create a comprehensive instruction dataset with 500K samples and three specific training objectives.
+During the training phase, we utilized only the training subsets of these benchmarks, while the inference phase involved evaluation using their respective test sets.
+3.3 Details of training and testing
+In this section, we detail the technical architecture of ChatReID.
+The ChatReID framework, illustrated in Figure 3, integrates the structure of Qwen2-VL [39] due to its Naive Dynamic Resolution architecture, which effectively addresses the challenges posed by variable image resolutions in person Re-ID tasks.
+In Stage 1, both the encoder and decoder are initialized with a pre-trained VLM to leverage existing language-vision knowledge.
+We conduct full-parameter training on both the vision encoder and language decoder of Qwen2-VL 2B to enhance the encoder’s focus on pedestrian features, while training the decoder ensures that captions accurately reflect pedestrian attributes.
+We standardize images with resolutions below 256$\times$128, resizing them to 256$\times$128 while leaving higher-resolution images unchanged.
+This adjustment leverages our experience with pedestrian tasks, and moderate image compression enhances training efficiency.
+In Stage 2, we carry over the encoder from Stage 1 but replace the decoder to continue full-parameter training.
 The previous decoder’s strong captioning focus could hinder outcomes in this stage, so we use a freshly pre-trained decoder instead.
@@
 Images larger than 384$\times$192 were rescaled to 384$\times$192, while images below this threshold were left unaltered.
-3.3 Decoder Post-training
-After compressing the general pedestrian discrimination information via the above two stages, ChatReID has foundational skills in person identity matching.
-During this stage, we fine-tune the language decoder to improve the practicality of GRP, deepening the model’s understanding of Re-ID objectives and further enhancing its instruction-following abilities for effective real-world deployment.
-3.3.1 Data Engine
-We rethink the assessment of the person Re-ID problem and formulate a question construction in the style of Visual Question Answering (VQA) utilizing a multi-modal large language model from a practical application perspective.
+In Stage 3, we inherit the full structure from Stage 2, training with the encoder frozen.
+With pedestrian recognition skills already developed, this stage focuses on enhancing the model’s instruction-following capabilities.
+The input image resolution is preserved without any adjustments.
+We rethink the assessment of the person Re-ID problem and formulate a question construction in the style of Visual Question Answering (VQA) utilizing a multi-modal language model from a practical application perspective.
 The instruction tuning data amounts to 500K.
@@
 The most common type of query involves providing a text or image query alongside a gallery set, with the goal of identifying images in the gallery that match the query’s identity.
-However, for current LVLMs, retrieving an indefinite-length sequence of images from an indefinite gallery set based on specific criteria is highly challenging.
+However, for current VLMs, retrieving an indefinite-length sequence of images from an indefinite gallery set based on specific criteria is highly challenging.
 To address this, we transform this complex retrieval task into a multi-turn best-choice problem through engineering enhancements.
@@
 Alternatively, users can start with image-based retrieval and add textual details to address complexities such as clothing changes or long-term recognition tasks.
-By integrating multi-modal information, this approach significantly improves both the flexibility and accuracy of pedestrian re-identification, advancing its real-world applicability.
+By integrating multi-modal information, this approach significantly improves both the flexibility and accuracy of person Re-ID, advancing its real-world applicability.
 To support this, we design a training objective for handling mixed queries with multi-modal inputs.
-3.3.2 Post-training
-We inherit the full structure from Stage 2, training with the encoder frozen.
-With pedestrian recognition skills already developed, this stage focuses on enhancing the model’s instruction-following capabilities.
-The input image resolution is preserved without any adjustments.
 4 Experiments
-4.1 Experimental Setups
+4.1 Experimental Setup
+Implementation details are provided in the supplementary.
 Datasets.
-We adhere to the datasets used in Stage 3, using only the test sets.
-We utilize the commonly used metrics in person Re-ID, such as Cumulative Match Characteristic (CMC) and Mean Average Precision (mAP).
-Results are presented in terms of mAP/Rank-1.
+We use the test sets from the datasets employed in Stage 3.
+The evaluation is conducted using standard metrics in person Re-ID, including Cumulative Match Characteristic (CMC) and Mean Average Precision (mAP).
+Results are reported in terms of mAP and Rank-1 accuracy.
 Evaluation strategy.
-Evaluating LVLMs for person Re-ID in a VQA format presents two main challenges.
-First, the large gallery sets in person Re-ID datasets exceed the token limits of current LVLMs.
-Second, retrieving an indefinite sequence of images based on specific criteria from a vast gallery is difficult for these models.
-To address these issues, we implement several engineering optimizations.
-We begin by training a base model to filter the gallery, calculating feature similarity between the query image and gallery images.
-Only images with similarity above a threshold $\tau$ are retained, significantly reducing gallery size and improving evaluation efficiency without affecting accuracy.
-Additionally, as described in Stage 3, we convert this retrieval task into a multi-turn best-choice problem.
-Finally, we sample non-overlapping images from the filtered gallery and pair each selected image with the query and rank them based on response confidence, producing the final similarity list.
+Evaluating VLMs for person Re-ID in a VQA format presents two primary challenges.
+First, the large gallery sets in person Re-ID datasets often exceed the token limits of current VLMs.
+Second, retrieving a sequence of images of variable length based on specific criteria from such vast galleries is inherently difficult for these models.
+To address these challenges, we implement several engineering optimizations.
+We first adopt a baseline model (e.g., ResNet-50) to filter the gallery by calculating feature similarity between the query image and gallery images.
+Only images with similarity above a threshold $\tau$ are retained, significantly reducing the gallery size while maintaining accuracy and improving evaluation efficiency.
+Next, we reformulate the retrieval task as a multi-turn best-choice problem.
+Finally, we sample non-overlapping images from the filtered gallery, pairing each selected image with the query and ranking them based on response confidence, ultimately producing the final similarity list.
 Table 1: Comparison of ChatReID with SOTA methods across three standard person Re-ID datasets.
@@
 ChatReID 96.4 97.2 87.5 90.1 91.3 92.7
-Table 3: Comparison of ChatReID with SOTA methods across two CC Re-ID datasets.
+Table 2: Comparison of ChatReID with SOTA methods across two cloth-changing Re-ID (CC Re-ID) datasets.
 Methods LTCC PRCC
@@
 As shown in Tab.
-2, ChatReID achieves impressive results across all standard person Re-ID datasets, particularly excelling on challenging datasets.
-The most notable improvement is observed on the complex and large-scale MSMT17 dataset, where ChatReID achieves a 15.1% increase in mAP.
-This substantial gain highlights the strength of our three-stage tuning strategy, which effectively extracts robust pedestrian-specific features, thereby improving model resilience in demanding scenarios.
-On the simpler Market1501 dataset, the performance gain is smaller, with a 2.9% increase in mAP.
-This modest improvement is expected, as existing methods have already achieved high accuracy on this dataset.
-Nonetheless, ChatReID’s measurable boost in performance demonstrates the consistent effectiveness of our approach, even when applied to datasets with strong baseline results.
-CC Re-ID.
+1, ChatReID achieves remarkable results across all standard person Re-ID datasets.
+The most notable improvement is observed on the challenging and large-scale MSMT17 dataset, where ChatReID achieves a 15.1% improvement in mAP.
+This substantial gain demonstrates the effectiveness of our three-stage tuning strategy, which effectively extracts robust, pedestrian-specific features, thereby improving the model’s resilience in demanding scenarios.
+Such performance improvement further validates the effectiveness of ChatReID.
+Cloth-changing Re-ID (CC Re-ID).
 The results in Tab.
 2 indicate that ChatReID demonstrates significant mAP enhancements over IRM on both the LTCC and PRCC datasets, with increases of 24.3% and 23.3% in mAP, respectively.
-Although both IRM and ChatReID incorporate several person Re-ID settings, ChatReID outperforms for two primary reasons.
-First, all three stages of our tuning are based on image-text pairs, where textual information aids the visual encoder in learning variations in clothing more effectively.
-Second, we introduced innovative attribute-based training tasks in Stage 2, which enhance the model’s ability to recognize cloth-changing scenarios.
-Table 4: Comparison of ChatReID with SOTA methods across two VI-ReID datasets.
+Based on our analysis, this improvement can be attributed to two main reasons.
+First, all three tuning stages are based on image-text pairs, where textual information aids the visual encoder learn variations in clothing more effectively.
+Second, we introduced innovative attribute-based training tasks in Stage 2, which enhance the model’s ability to recognize CC scenarios.
+Table 3: Comparison of ChatReID with SOTA methods across two visible-infrared person ReID (VI-ReID) datasets.
 Methods SYSU-MM01 RegDB
@@
 ChatReID 83.6 86.8 95.8 96.5
-Table 5: Ablation Study. Performance comparison of different stage combinations in our hierarchical progressive learning strategy across ten benchmarks.
+Table 4: Ablation Study.
+Performance comparison of different stage combinations in our hierarchical progressive learning strategy across ten benchmarks.
+The abbreviation ‘MM01‘ refers to the SYSU-MM01 dataset, ‘CUHK‘ refers to the CUHK-PEDES dataset, and ‘ICFG‘ refers to the ICFG-PEDES dataset.
 Methods Standard ReID CC-ReID VI-ReID T2I-ReID Avg.
-Market1501 MSMT17 CUHK03 PRCC LTCC SYSU-MM01 RegDB CUHK-PEDES. ICFG-PEDES RSTPReid
+Market1501 MSMT17 CUHK03 PRCC LTCC MM01 RegDB CUHK ICFG RSTPReid
 Stage 3 59.2 41.5 60.3 31.2 35.4 11.4 12.7 65.1 49.9 43.5 41.0
@@
 ChatReID 80.1 83.8 70.5 72.9 73.1 75.0
-VI-ReID.
+Visible-infrared person ReID (VI-ReID).
 As shown in Tab.
-4, ChatReID also demonstrates significant improvements on VI-ReID datasets.
+3, ChatReID also demonstrates significant improvements on VI-ReID datasets.
 Specifically, it achieves an 11.8% mAP increase on SYSU-MM01 and a 10.7% mAP increase on RegDB.
-These gains are primarily due to our approach in Stage 2, where we treated the VI-ReID task as an image-image matching task and slightly increased the proportion of VI data in the tuning process.
-This adjustment allowed the model to learn VI matching more effectively, resulting in enhanced performance in cross-modality Re-ID tasks.
+These gains are primarily attributed to our approach in Stage 2, where we treated the VI-ReID task as an image-image matching task and slightly increased the proportion of VI data in the tuning process.
 Text-to-Image Person Re-ID.
 As shown in Tab.
-6, ChatReID achieves remarkable gains in text-to-image person Re-ID, especially on the ICFG-PEDES and RSTPReid datasets, where prior performance was relatively low.
+5, ChatReID achieves remarkable gains in text-to-image person Re-ID, especially on the ICFG-PEDES and RSTPReid datasets, where previous performance was relatively low.
 ChatReID reaches an mAP of 70.5% on ICFG-PEDES, marking a 25.6% improvement, and an mAP of 73.1% on RSTPReid, a 20.2% increase.
-These significant gains can be attributed to two key factors.
-First, existing LVLMs already possess a baseline level of text-to-image matching capability, as evidenced in our ablation studies.
-Second, throughout all three training stages, we utilize a dataset structure based on image-text pairs, which substantially enhances the model’s ability to learn pedestrian-specific text-to-image matching.
+These gains are attributed to two factors: existing VLMs have a baseline text-to-image matching capability, and our image-text pair dataset structure enhances person-specific matching across all training stages.
 Attribute-Based Person Retrieval.
-To our knowledge, there is currently no straightforward and effective method for the Attribute-Based Person Retrieval task.
+To the best of our knowledge, there is currently no straightforward and effective method for the Attribute-Based Person Retrieval task.
 Existing text-to-image person Re-ID methods primarily focus on retrieving images based on complete textual descriptions rather than specific attributes.
-To address this gap, we introduced attribute-based retrieval as a training objective in Stage 3 of our model.
+To address this gap, we introduced attribute-based retrieval as a training objective in Stage 3.
 To assess the effectiveness of our approach, we reorganized the dataset by randomly sampling $N$ images from an attribute-annotated dataset.
 Our goal was to retrieve images that matched specified attributes, such as clothing or accessories.
-Experimental results indicate that ChatReID can efficiently retrieve images based on attribute criteria, underscoring its practical value for real-world applications.
+Experimental results indicate that ChatReID can efficiently retrieve images based on attribute criteria, highlighting its practical value for real-world applications.
 Additional experimental results are presented in the supplementary material.
 Adaptive Multi-Modal Person Retrieval.
-Considering practical application needs, we design a training task for Adaptive Multi-Modal Person Retrieval.
+Considering practical application requirements, we design a training task for Adaptive Multi-Modal Person Retrieval.
 To evaluate the model’s performance on this task, we conducted experiments that included additional textual information as auxiliary input for image-image retrieval tasks.
-The experimental results demonstrate that ChatReID can efficiently handle person retrieval with multi-modal inputs and that incorporating multi-modal information enhances performance in person Re-ID task.
+The experimental results demonstrate that ChatReID can efficiently handle person retrieval with multi-modal inputs.
 Detailed results and analyses are provided in the supplementary material.
 4.3 Ablation Study
-To validate the effectiveness of the three-stage hierarchical progressive tuning strategy, we compare experimental results from different stage combinations.
+Three-Stage HPT Strategy.
+To validate the effectiveness of our three-stage HPT strategy, we compare experimental results across different stage combinations, as shown in Tab.
+4.
+Initially, we report the results from directly applying Stage 3 training.
+While Stage 3 achieves baseline performance, the results are suboptimal, indicating that current VLMs have limited inherent capabilities for image-to-image and image-text pedestrian matching and require specialized training for optimal performance.
+Our experiments further reveal that combining only Stage 1 and Stage 3 made it difficult for the model to converge, and thus, those results were not reported.
+However, incorporating Stage 2 into the training process significantly improved the model’s performance, achieving an average mAP of 73.5%.
+The fine-grained image retrieval tuning in Stage 2 led to substantial improvements, providing strong evidence for the effectiveness of our approach.
+Finally, when Stage 1 training is also included, the model’s performance is further enhanced, reaching state-of-the-art (SOTA) results across ten benchmarks.
+This final improvement underscores the overall effectiveness of our hierarchical three-stage tuning strategy.
+VLMs Model Size and Performance Trade-off.
+In our validation experiments, we evaluated VLMs of different sizes under each setting.
 The results are shown in Tab.
-5.
-We first report results from the application of Stage 3 training directly to the pre-trained Qwen2-VL model.
-The results show that while the pre-trained Qwen2-VL achieves some baseline performance, the results are suboptimal, indicating that current LVLMs possess limited inherent image-image and image-text pedestrian matching capabilities but require specialized training for optimal performance.
-Performance on the VI-ReID task was particularly low, likely due to the limited proportion of low-illumination data in the training set, which is insufficient for training a robust model.
-In our experiments, we found that combining only Stage 1 and Stage 3 made it difficult for the model to converge; thus, we did not report those results.
-However, adding Stage 2 significantly improved model performance, achieving an average mAP of 73.5%, which demonstrates practical utility.
-This validates the effectiveness of our Stage 2 training design.
-When Stage 1 training is also included, the model’s performance further improves, reaching SOTA results across ten benchmarks.
-This final enhancement reinforces the overall effectiveness of our hierarchical three-stage tuning strategy.
-5 Conclusion
-This paper introduces a novel framework for person re-identification (Re-ID), ChatReID, that leverages the advanced capabilities of large vision-language models (LVLMs).
-Unlike traditional methods, which are limited by rigid templates and lack of adaptability, our system enables freeform, user-driven queries, offering greater flexibility and natural interaction.
-By implementing a Hierarchical Progressive Tuning (HPT) strategy, we progressively enhance the model’s ability to achieve fine-grained identity-level retrieval.
-Our approach, combining Incremental Pre-training, Multi-task Joint Training, and Decoder Post-training , delivers SOTA performance across various Re-ID benchmarks.
-The system’s VQA inference format simplifies complex Re-ID tasks, making it accessible to non-experts, while its flexibility allows dynamic input combinations and adjustments.
-Our work establishes a strong foundation for future advancements in person Re-ID, showcasing the potential of LVLMs in practical applications.
+6.
+As the model size increases, accuracy improves.
+Both the 2B and 7B models achieve comparable performance.
+This can be attributed to the fact that the Re-ID task is relatively less complex compared to VLM training scenarios, allowing the 2B model to effectively capture the necessary representations.
+Given that the 7B model incurs substantially higher computational costs without yielding significant performance gains, we chose the 2B model.
+Table 6: Performance on VLMs with different model size. The data in the table represents Rank-1.
+#Parameter MSMT17 LTCC RegDB CUHK-PEDES
+0.5 B 38.1 27.4 4.3 55.1
+2 B 90.1 82.7 96.5 83.8
+7 B 91.3 82.5 96.3 87.7
+5 Broad Impact and Conclusion
+Broad Impact.
+We utilize publicly available, high-quality academic datasets for our research.
+These datasets were curated by various institutions across diverse scenarios, which helps mitigate potential biases in the images.
+It is important to note that, due to privacy policies, we do not use the DukeMTMC [31] dataset.
+We will release the code, datasets, and models upon acceptance to facilitate further research.
+To ensure responsible use, we will adhere to a strict application protocol to prevent our work from being used for any unethical or illegal purposes.
+Conclusion.
+This paper introduces ChatReID, a novel framework for person re-identification (Re-ID) that leverages the advanced capabilities of vision-language models (VLMs).
+Our ChatReID is a text-side-dominated framework for person Re-ID.
+Different from traditional methods that calculates image feature distance as similarity, ChatReID interprets the task requirements from the input text description and performs similarity inference on the given person images accordingly.
+By implementing a Hierarchical Progressive Tuning (HPT) strategy, we progressively enhance the model’s ability to achieve fine-grained, identity-level retrieval.
+The system’s VQA-based inference format simplifies complex Re-ID tasks, making it more accessible to non-experts, while its flexibility allows dynamic input combinations and adjustments.
+Our work lays a strong foundation for future advancements in person Re-ID, showcasing the potential of LVLMs in practical applications.